OpenStudy (anonymous):

x/5 + 9 = 4 A. 65 B. -25 C. 5 D.20

OpenStudy (rushwr):

\[\frac{ x }{ 5 } +9 = 4\] \[\frac{ x }{ 5 } = 4 -9\] \[\frac{ x }{ 5 } = -5\] \[x= -5 * 5 = -25\]

OpenStudy (rushwr):

So the answer would be the second one !

OpenStudy (anonymous):

Thank you so much ^_^

OpenStudy (anonymous):

-25